After leaving Dallas, Parker drove 300 miles at x mph and then decreased his speed by 10 mph and drove an additional 400 miles. Make a table with the given information and write an expression for his total traveling time. Evaluate the expression for x = 60.
Answer by xcentaur(357)

DataFirst caseSecond case
distance300 miles400 miles
speedx mph(x-10)mph
time=distance/speed300/x400/(x-10)

total travelling time=(300/x)+(400/(x-10))
for x=60,total travelling time is:
=(300/x)+(400/(x-10))
=(300/60)+(400/(60-10))
=(300/60)+(400/50)
=5+8
=13 hours


Hope this helps,
Prabhat